# What constitutes a Deep Fake? The blurry line between legitimate processing and manipulation under the EU AI Act

### v1

@@
Abstract
When does a digital image resemble reality? The relevance of this question increases as the generation of synthetic images—so called deep fakes—becomes increasingly popular. Deep fakes have gained much attention for a number of reasons—among others, due to their potential to disrupt the political climate. In order to mitigate these threats, the EU AI Act implements specific transparency regulations for generating synthetic content or manipulating existing content. However, the distinction between real and synthetic images is— even from a computer vision perspective—far from trivial. We argue that the current definition of deep fakes in the AI act and the corresponding obligations are not sufficiently specified to tackle the challenges posed by deep fakes. By analyzing the life cycle of a digital photo from the camera sensor to the digital editing features, we find that: (1.) Deep fakes are ill-defined in the EU AI Act. The definition leaves too much scope for what a deep fake is. (2.) It is unclear how editing functions like Google’s “best take” feature can be considered as an exception to transparency obligations. (3.) The exception for substantially edited images raises questions about what constitutes substantial editing of content and whether or not this editing must be perceptible by a natural person.
@@ Abstract
1 Introduction
@@
Generative AI tools that can generate photorealistic images, or manipulate existing photos in a convincing manner, are now widely available. The term “deep fake” has been coined, referring to images generated or manipulated using deep neural networks. Actual deep fakes, and allegations of their use, play an increasingly important role in political debates and election campaigns as well as in pornographic content [Becker(2024)]. Given the potential impact of manipulated images, regulators increasingly identify a need to regulate the use of such images. Countries that have already initiated efforts to regulate deep fakes include the US, China, the UK, and South Africa [Ramluckan(2024), Murray(2024)].
@@
The European Union (EU)’s AI Act includes a transparency rule for AI systems generating deep fakes and other synthetic content. The AI Act’s definition relates to “manipulated” content, which raises the question of how to differentiate between manipulation and permissible editing. Photos have never been an objective depiction of reality. In the context of photojournalism, an author has called for framing “the production and use of reality images as ‘mediated communication’ rather than an ‘objective truth’.” [Newton(2013)] The photographer’s location and the focal length used, the depth of field and the lighting can create completely different visual impressions of the same object, even in analog photography and without post-processing. As seen in advertising photography, this is often done with manipulative intent. Digital post-processing opens up additional possibilities, such as adaptations of white balance and the color palette or the removal of spots caused by dust on the image sensor. In recent years, an increasing range of AI-based post-processing tools has become available to amateur photographers and smartphone users. The camera app of some Samsung phones automatically replaces depictions of the moon by an improved and adapted image (see Fig. 1). Google’s recently presented “Best Take” feature allows swapping faces, enabling users to ensure that all persons depicted in a group photo are shown with a smile and their eyes open. This paper aims to open the discussion on where to draw the line between legitimate processing and manipulative deep fakes. It focuses on the interpretation of the European AI Act, but also raises fundamental questions.
@@
While previous work has discussed how the AI Act impacts image forensics in regard to deep fake detection [Lorch et al.(2022)], we focus on the generation and manipulation of content with deep fakes and the corresponding transparency regulations. The literature has already discussed whether AI systems which generate deep fakes can be considered as high-risk systems (Article 6) [^1] Legal citations without reference to a statute always refer to the EU AI Act (Regulation (EU) 2024/1689)). or general-purpose AI systems (Article 3(66)) [Block(2024)]. We do not include this discussion in this paper.
@@
The remainder of this paper is structured as follows: Section 2 presents common image processing techniques, both traditional and with the support of AI. In Section 3, we provide the background on the regulation of deep fakes in the European AI Act. Based on the two previous sections, we discuss the challenges and the edge cases of deep fake regulation in Section 4, and conclude with an outlook in Section 5.
@@
Figure 1: Details of moon photographs taken with a Samsung Galaxy S20 5G smartphone by Thomas Wagner.
2 Image processing – Lifecycle of a digital photo
2.1 Traditional
In order to be intelligible to a viewer, digital photos always have to be processed. The following steps are often performed automatically (e.g., within the Android camera subsystem), and without the user necessarily being aware of them [Minhaz(2020)]:
@@ 2.1 Traditional
Figure 2: Demonstration of dust removal (detail of a photo, dust removal performed with Luminar Neo without human intervention). Image best viewed in colour.
@@
  • The image contrast around edges may be increased so the image appears sharper.
Instead of performing these steps in the camera, professional photographers often choose to work on a mostly uncorrected (raw) image to get more control over the process. For instance, they may want to correct distortions caused by the camera perspective in addition to those introduced by the lens, or adapt colors as an artistic choice instead of just correcting for lighting. Numerous additional filters and editing tools are available and commonly used. These include the removal of spots caused by dust on the image sensor (a common issue on cameras with interchangeable lenses), as well as sharpening and softening effects.
@@
All these processing steps influence the perception of an image by a human viewer, but are common or even necessary.
2.2 AI-supported post-processing
Photographers use AI-based post-processing to an increasing extent: While marketing claims about AI use are not a reliable indicator for a tool to be classified as AI, available tools provide functions that cannot be plausibly achieved without the use of AI.
Some of these functions replace processing that was previously available, but required a greater extent of human intervention. The removal of image sensor dust commonly required users to mark the spots to be removed, but can now be performed automatically (see Fig. 2). Similarly, power lines can be easily removed from an image as well. [^2] See, for instance, https://skylum.com/en/how-to/how-to-remove-dust-spots-power-lines-automatically. In order to blur the background in post-processing (which achieves a similar effect to choosing a wide aperture during the photo shoot, see Fig. 3), photographers previously needed to mark the foreground; nowadays, this task can often be performed without intervention.
Other functions are only realistically possible due to the use of AI. Adobe Photoshop includes an “AI Image Extender” feature that allows to “magically add more background” [^3] https://www.adobe.com/products/photoshop/generative-expand.html# using generative AI. Luminar Neo’s BodyAI tool features a slider to “slim or add volume to a subject’s torso” [^4] https://userguide.skylum.com/hc/en-us/articles/11542260006930-Working-with-the-Portrait-Tools.
AI-based post-processing is commonly available on modern smartphones as well. This makes photo editing options easily accessible to many users, and they might not even be aware of the functionality.
@@ 2.2 AI-supported post-processing
Figure 3: Demonstration of background blurring with Luminar Neo’s “Portrait Bokeh” tool, showing details of an image
@@
Google has implemented a feature called “Best Take” on Pixel smartphones, which is focused on improving pictures of persons (individuals and groups). [^7] https://store.google.com/intl/en/ideas/articles/pixel-best-take/ The core idea of that feature has been described in a research paper [Shah and Kwatra(2012)]. A series of photos is taken instead of a single one. The software then analyzes the facial expressions and selects the best image. Moreover, in a group photo, individual faces are seamlessly replaced by better versions (e.g., from a different take of the same photo series in which the respective person’s eyes are open).
@@
3 Regulation of deep fakes in the European AI Act
Article 3(60) defines deep fakes, and Article 50 states resulting transparency obligations. Recitals 132 to 137 include further information on the interpretation of the transparency obligations. In the following sections, we discuss the core definition of deep fakes before presenting the resulting obligations.
3.1 Definition of deep fakes in the AI Act
According to Article 3(60),
@@ 3.1 Definition of deep fakes in the AI Act
According to the definition, four conditions need to be met in order for a content to be considered a deep fake. First the content needs to be generated or manipulated. Generation can be understood as prompt based content creation [Gils(2024)]. In contrast, manipulation refers to given input data, for example, an already taken image, being altered. As discussed in section 2, an image taken by a camera never shows the (true) reality but instead is a result of the image processing chain as well as settings chosen by the photographer. For image content, it is problematic to define the difference between traditional image processing steps and novel deep fake manipulation. We get back to this point in section 4.1.
@@
Second, image, audio, or video content is needed for a deep fake. Other content, such as text, are not considered deep fakes in the AI Act  [^8] There are, however, transparency obligations in Article 50 which also covers synthetic text content. We discuss this point in section 3.2.
@@
Third, the content needs to resemble existing objects, places, entities, or events to be considered a deep fake. Thus, a link to the real world is necessary. Unicorns or monsters, which only give a photorealistic impression, are not considered deep fakes [Becker(2024)]. Recital 134 includes a slightly modified version of the definition of a deep fake:
@@
[…]AI system to generate or manipulate image, audio or video content that appreciably resembles existing persons, objects, places, entities or events and would falsely appear to a person to be authentic or truthful (deep fakes)[…]
@@
Adding the word appreciably might slightly change the meaning of the definition, emphasizing the resemblance to existing objects [Gils(2024)]. The question arises “whether the requirement of resemblance demands identicality or similarity to the existing subjects.” [Gils et al.(2024)]. For example, in case of a face the question is whether the face has to belong to a real person or whether the face has to be sufficiently similar to real persons. In this context, it has been argued that the difference between the definitions only shows the lack of regulatory precision and has no practical influence [Łabuz(2024)].
@@
Finally, the content must appear to a person to be authentic or truthful. Thus, the impression to a natural person is important. Generated or manipulated content (such as adversarial perturbed images [Szegedy(2013)]) that is only classified as an existing person by an algorithm is not considered a deep fake. It has been argued that the requirement of an impression to a person compared to an average person broadens the scope considerably [Block(2024)]. For example, some images might be perceived as deep fakes by children but not by adults [Block(2024)]. It is unclear, which background knowledge and abilities of a person must be taken into consideration when defining content as deep fake. Moreover, available editing options in modern image processing software challenge the definition of a deep fake, see section 4.1.
@@
Not only the EU AI act includes a regulation of deep fakes. In other EU legislative acts, deep fakes are also defined [Łabuz(2023)]. Specifically the Digital Services Act (DSA) contains a similar definition in Article 35(1)(k) DSA, which might pose challenges for the interplay between DSA and AI act. [^9] See the cited paper for more information. However, while previous versions of the AI act had different definitions of deep fakes, now at least a basic level of consistency between the AI act and other regulations such as the DSA has been reached [Łabuz(2024)]. This basic level of consistency allows partially consistent obligations between the different legislative acts.
3.2 Transparency Obligations
@@
Article 50 of the AI Act includes the main transparency regulations and obligations concerning synthetic content and deep fakes. These obligations apply regardless of whether the underlying system is a high-risk system or general-purpose AI, as noted in recital 132.
Article 50 (2) demands that
@@
The first observation is that Article 50(2) only covers the providers (see Article 3(3) for a definition of providers). In contrast, deployers (see Article 3(4)) are governed by Article 50(4).
@@
Second, Article 50(2) mentions only the generation of synthetic content. Here the question arises whether the manipulation of existing content is also covered by the provision in Article 50 (2) [Gils(2024)]. There are strong reasons to believe it is. In the latter part of the first sentence, Article 50(2) mentions the obligation to mark the synthetic content as generated or manipulated. Recital 133 explicitly also mentions the manipulation of existing inputs. Thus, we assume here that the generation and manipulation of input are covered by Article 50(2).
@@
In contrast to the core deep fake regulation in Article 50(4), here, text content is also listed. Thus the obligation of Article 50(2) include deep fakes but also extends beyond core deep fake regulation.
@@
The main obligation is to mark the synthetic content in a machine-readable format to ensure that the content is detectable as artificially generated or manipulated. On a technical side, watermarking synthetic outputs is far from a trivial task [Gils(2024)], although Recital 133 explicitly mentions some techniques. It remains to be seen whether the community will invest some joint effort to develop global standards [Romero Moreno(2024)].
@@
Article 50(2) does not explicitly mention that the marking of the synthetic output has to be detectable by human viewers [^10] We specifically cover image content in our paper. per se. Thus, it remains unclear whether there has to be a human-visible declaration of the synthetic content [Gils(2024)]. However, Article 50(5) states that the information referred to in Article 50(2) has to be provided to natural persons in a clear and distinguishable manner. Marking content only in a machine-readable way without visibility to humans would be deceptive. So in practice, it does not matter whether Article 50(2) alone or in conjunction with Article 50 (5) requires some human-visible marking. Additionally providers of (large) social media platforms already have an obligation to label deep fakes through Article 35(1) DSA [Block(2024)].
@@
In contrast to the aforementioned minor issues in the transparency obligation, two exceptions in Article 50(2) pose a major difficulty. Article 50(2) states that
@@
This obligation [of Article 50(2)] shall not apply to the extent the AI systems perform an assistive function for standard editing or do not substantially alter the input data provided by the deployer[…]
@@
So the first exception targets systems that perform assistive functions for standard editing tasks. However, the legislator’s intention remains unclear [Block(2024)]. There is no legal definition nor a clarification in the recitals for either of the terms “assistive function” and “standard editing”. What characterises standard editing given the existence of modern AI based image editing tools, such as Samsung’s scene optimizer or Google’s Best take? The threshold for an assistive function or a standard editing tool is not fixed.
@@
From our point of view, both exceptions [^11] There is also a third exception for law enforcement purposes which we do not consider in our paper pose a serious challenge for existing image editing tools on smartphones and computers. We discuss this issue in depth in section 4
The other main transparency obligation regarding deep fakes is given in Article 50(4) subparagraph 1. Here deep fakes are specifically mentioned:
@@
In contrast to Article 50(2), this obligation only targets the deployers of an AI system. Furthermore, only deployers of systems processing image, audio, or video are regulated. Text systems, on the other hand, are not considered as deep fakes under the AI Act, and are regulated in Article 50(4) subparagraph 2. It is unclear whether Article 50(4) subparagraph 1 requires the AI content itself to be labeled—or whether labelling the system, for example through a banner on the webpage, is sufficient. However, Recital 134 clearly indicates that the AI content itself needs a label disclosing the artificial content. Thus, deployers of an AI system that create deep fakes need to ensure that humans can observe the origin [Romero Moreno(2024)].
@@
Compared to the non-trivial exceptions in Article 50(2), the exceptions for the deployers in Article 50(4) are clearer. There are exceptions and limitations to this obligation for the law enforcement regime as well as for evidently artistic, creative, satirical, fictional or analogous work. The latter exception refers to the freedom of expression and artistic creation protected by Articles 11 and 13 EU Charter [Romero Moreno(2024)].
@@
The interplay between the obligations for the provider in Article 50(2) and the obligation for the deployer in Article 50(4) is complex, and remains largely unclear. Recital 134 explicitly mentions that further to the technical solutions employed by the providers of the AI system, the deployers also have to comply with the transparency obligations. It seems questionable if double labelling adds a benefit for the transparency. Standardized labeling efforts are needed [Gils(2024)].
@@
Both obligations try to regulate a technology that is, without the underlying intention of the user, neutral [Łabuz(2024), De Ruiter(2021)]. However, the overarching theme of both obligations is to ensure that even in cases where AI systems can create large amounts of synthetic images, humans are able to differentiate between synthetic and real images, see also Recital 133. This focus on the recipients of the synthetic content has already been criticised since it does not focus on the victims of deep fakes, for example in the case of (deep) pornographic images or videos [Veale and Borgesius(2021), Łabuz(2024), Meskys et al.(2020)], which can cause severe harm to people. However, the non-compliance of providers or deployers of (commercial) AI systems with Article 50 can be subject to a fine of up to €15,000,000, or, if the offender is an undertaking, up to 3% of its total worldwide annual turnover for the preceding financial year, whichever is higher (see Article 99). Thus, non-compliance can be very costly for a company.
3.3 Summary of main critical legislative issues in Article 50 AI act

### v2

@@
Abstract.
When does a digital image resemble reality? The relevance of this question increases as the generation of synthetic images—so-called deep fakes—becomes increasingly popular. Deep fakes have gained much attention for a number of reasons—among others, due to their potential to disrupt the political climate. In order to mitigate these threats, the EU AI Act implements specific transparency regulations for generating synthetic content or manipulating existing content. However, the distinction between real and synthetic images is—even from a computer vision perspective—far from trivial. We argue that the current definition of deep fakes in the AI Act and the corresponding obligations are not sufficiently specified to tackle the challenges posed by deep fakes. By analyzing the life cycle of a digital photo from the camera sensor to the digital editing features, we find that: (1.) Deep fakes are ill-defined in the EU AI Act. The definition leaves too much scope for what a deep fake is. (2.) It is unclear how editing functions like Google’s “best take” feature can be considered as an exception to transparency obligations. (3.) The exception for substantially edited images raises questions about what constitutes substantial editing of content and whether or not this editing must be perceptible by a natural person.
@@ Abstract.
Deep Fakes, EU AI Act, Image Processing, Legal Aspects, Transparency Regulations
journalyear: 2025conference: Symposium on Computer Science and Law; March 25–27, 2025; München, Germanybooktitle: Symposium on Computer Science and Law (CSLAW ’25), March 25–27, 2025, München, Germanydoi: 10.1145/3709025.3712218isbn: 979-8-4007-1421-4/25/03
1. Introduction
@@
Generative AI tools that can generate photorealistic images or manipulate existing photos in a convincing manner are now widely available. The term “deep fake” has been coined, referring to images generated or manipulated using deep neural networks. Actual deep fakes and allegations of their use play an increasingly important role in political debates and election campaigns as well as in pornographic content (Becker, 2024; Kristof, 2024). Given the potential impact of manipulated images, regulators increasingly identify a need to regulate the use of such images. Countries that have already started initiatives to regulate deep fakes include the US, China, the UK, and South Africa (Ramluckan, 2024; Murray, 2024).
@@
The European Union (EU)’s AI Act includes a transparency obligation for AI systems generating deep fakes and other synthetic content. The AI Act’s definition relates to “manipulated” content, which raises the question of how to differentiate between manipulation and permissible editing. Photos have never been an objective depiction of reality. In the context of photojournalism, an author has called for framing “the production and use of reality images as ‘mediated communication’ rather than an ‘objective truth”’ (Newton, 2013). The photographer’s location and the focal length used, the depth of field and the lighting can create completely different visual impressions of the same object, even in analog photography and without post-processing. As seen in advertising photography, this is often done with manipulative intent. Digital post-processing opens up additional possibilities, such as adaptations of white balance and the color palette or the removal of spots caused by dust on the image sensor. In recent years, an increasing range of AI-based post-processing tools has become available to amateur photographers and smartphone users. The camera app pre-installed on some Samsung phones automatically replaces depictions of the moon by an improved and adapted image (see Fig. 1). Google’s recently presented “Best Take” feature allows swapping faces, enabling users to ensure that all persons depicted in a group photo are shown with a smile and their eyes open. This paper aims to open the discussion on where to draw the line between legitimate processing and manipulative deep fakes. It focuses on the interpretation of the European AI Act, but also raises fundamental questions.
@@
While previous work has discussed how the AI Act impacts image forensics in regard to deep fake detection (Lorch et al., 2022), we focus on the generation of deep fakes either from scratch or through the manipulation of (existing/real) content, and the corresponding transparency regulations. The literature has already examined whether AI systems that generate deep fakes can be considered high-risk systems (Article 6 AI Act [^1] Regulation (EU) 2024/1689. (AIA)) or general-purpose AI models (Article 3(66) AIA)) (Block, 2024). We do not include this discussion in this paper.
@@
The remainder of this paper is structured as follows: Section 2 presents common image processing techniques, both traditional and with the support of AI. In Section 3, we provide the background on the regulation of deep fakes in the European AI Act. Based on the two previous sections, we discuss the challenges and edge cases of deep fake regulation in Section 4, and conclude with an outlook in Section 6.
@@
[…]
An image of the full moon with significantly more detail than the left one.
Figure 1. Details of moon photographs taken with a Samsung Galaxy S20 5G smartphone by Thomas Wagner.
2. Image processing – Lifecycle of a digital photo
2.1. Traditional
Digital photos must always be processed to be intelligible to a viewer. The following steps are often performed automatically (e.g., within the Android camera subsystem), and without the user necessarily being aware of them (Minhaz, 2020):
@@ 2.1. Traditional
[…]
Parts of a building and a blue, cloudless sky with some darker spots (specks of dust).
@@
[…]
Parts of a building and a blue, cloudless sky.
Figure 2. Demonstration of dust removal (detail of a photo, dust removal performed with Luminar Neo without human intervention). Image best viewed in colour.
@@
  • The image contrast around edges may be increased, so the image appears sharper.
Instead of performing these steps with the camera, professional photographers often choose to work on a mostly uncorrected (raw) image to get more control over the process. For instance, they may want to correct distortions caused by the camera perspective in addition to those introduced by the lens or adapt colors as an artistic choice instead of just correcting for lighting. Numerous additional filters and editing tools are available and commonly used. These include the removal of spots caused by dust on the image sensor (a common issue on cameras with interchangeable lenses), as well as sharpening and softening effects.
@@
All these processing steps influence the perception of an image by a human viewer but are common or even necessary.
2.2. AI-supported post-processing
Photographers use AI-based post-processing to an increasing extent: Although marketing claims about AI use are not a reliable indicator for a tool to be classified as AI, available tools provide functions that cannot be plausibly achieved without the use of AI.
Some of these functions replace processing that was previously available but required human intervention to a greater extent. The removal of image sensor dust commonly required users to mark the spots to be removed, but can now be performed automatically (see Fig. 2). Similarly, power lines can be easily removed from an image as well [^2] See, for instance, https://skylum.com/en/how-to/how-to-remove-dust-spots-power-lines-automatically.. In order to blur the background in post-processing (which achieves a similar effect to choosing a wide aperture during the photo shoot, see Fig. 3), photographers previously needed to mark the foreground; nowadays, this task can often be performed without intervention.
Other functions are only realistically possible due to the use of AI. Adobe Photoshop includes an “AI Image Extender” feature that allows to “magically add more background” [^3] https://www.adobe.com/products/photoshop/generative-expand.html. using generative AI. Luminar Neo’s BodyAI tool features a slider to “slim or add volume to a subject’s torso” [^4] https://userguide.skylum.com/hc/en-us/articles/11542260006930-Working-with-the-Portrait-Tools..
AI-based post-processing is commonly available on modern smartphones as well. Thus, photo-editing options are easily accessible to many users, and they might not even be aware of the functionality.
@@ 2.2. AI-supported post-processing
[…]
A partial face in the foreground and a plant in the background. Both the face and the plant are sharply in focus.
@@
[…]
A partial face in the foreground and a plant in the background. The face is sharply in focus, while the plant is blurry.
@@
[…]
A partial face in the foreground and a plant in the background. The face is sharply in focus, while the plant is blurry.
Figure 3. Demonstration of background blurring with Luminar Neo’s “Portrait Bokeh” tool, showing details of an image
@@
Google has implemented a feature called “Best Take” on Pixel smartphones, which is focused on improving pictures of persons (individuals and groups). [^7] https://store.google.com/intl/en/ideas/articles/pixel-best-take/ The core idea of that feature has been described in a research paper (Shah and Kwatra, 2012). A series of photos is taken instead of a single one. The software then analyzes facial expressions and selects the best image. Moreover, in a group photo, individual faces are seamlessly replaced by better versions (e.g., from a different take of the same photo series in which the respective person’s eyes are open).
@@
3. Regulation of deep fakes in the European AI Act
Article 3(60) AIA defines deep fakes, and Article 50 AIA states resulting transparency obligations. Recitals 132 AIA to 137 AIA include further information on the interpretation of these transparency obligations. In the following sections, we discuss the core definition of deep fakes before presenting the resulting obligations.
3.1. Definition of deep fakes in the AI Act
According to Article 3(60) AIA,
@@ 3.1. Definition of deep fakes in the AI Act
According to the definition, four conditions need to be met in order for content to be considered a deep fake. First, the content needs to be generated or manipulated. Generation can be understood as (prompt-based) content creation (Gils, 2024). In contrast, manipulation refers to given input data—for example, an already taken image—being altered. As discussed in Section 2, an image taken by a camera never shows the (true) reality but instead is a result of the image processing chain as well as settings chosen by the photographer. For image content, defining the difference between traditional image processing steps and novel deep fake manipulation is problematic. We get back to this point in Section 4.1.
@@
Second, image, audio, or video content is needed for a deep fake. Other content, such as text, is not considered deep fakes in the AI Act [^8] There are, however, transparency obligations in Article 50 AIA, which also covers synthetic text content. We discuss this point in Section 3.2..
@@
Third, the content needs to resemble existing objects, places, entities, or events to be considered a deep fake. Thus, a link to the real world is necessary. Unicorns or monsters, which only give a photorealistic impression, are not considered deep fakes (Becker, 2024). Recital 134 AIA includes a slightly modified version of the definition of a deep fake:
@@
[…] AI system to generate or manipulate image, audio, or video content that appreciably resembles existing persons, objects, places, entities, or events and would falsely appear to a person to be authentic or truthful (deep fakes)[…]
@@
Adding the word appreciably might slightly change the meaning of the definition, emphasizing the resemblance to existing objects (Gils, 2024). The question arises “whether the requirement of resemblance demands identicality or similarity to the existing subjects.” (Gils et al., 2024). For example, in the case of a face, the question is whether the face has to belong to a real person or whether the face has to be sufficiently similar to real persons. In this context, it has been argued that the difference between the definitions only shows the lack of regulatory precision and has no practical influence (Łabuz, 2024).
@@
Finally, the content must appear to a person to be authentic or truthful. Thus, the impression to a natural person is important. Generated or manipulated content (such as adversarial perturbed images (Szegedy, 2013)) that is only classified as an existing person by an algorithm is not considered a deep fake. It has been argued that the requirement of an impression to a person compared to an average person considerably broadens the scope (Block, 2024). For example, some images might be perceived as authentic by children but not by adults (Block, 2024). It is unclear which background knowledge and abilities of a person must be taken into consideration when classifying content as a deep fake. Moreover, available editing options in modern image processing software challenge the definition of a deep fake, see Section 4.1.
@@
Deep fakes are not only regulated in the EU AI Act, but also in other legislative acts of the EU (Łabuz, 2023). Specifically, Article 35(1)(k) of the Digital Services Act (DSA) contains a similar wording to the AI Act’s definition of deep fakes, which might pose challenges for the interplay between DSA and AI act. [^9] See the cited paper for more information. However, while previous versions of the AI act had different definitions of deep fakes, now at least a basic level of consistency between the AI act and other regulations such as the DSA has been reached (Łabuz, 2024). This basic level of consistency allows for partially consistent obligations between the different legislative acts.
3.2. Transparency Obligations
@@
Article 50 AIA of the AI Act includes the main transparency regulations and obligations concerning synthetic content and deep fakes. These obligations apply regardless of whether the underlying system is a high-risk system or general-purpose AI, as noted in recital 132.
Article 50(2) AIA demands that
@@
The first observation is that Article 50(2) AIA only covers providers (see Article 3(3) AIA for a definition of providers). In contrast, deployers (see Article 3(4) AIA)) are governed by Article 50(4)AIA.
@@
Second, Article 50(2) AIA mentions only the generation of synthetic content. Here, the question arises whether the manipulation of existing content is also covered by the provision in Article 50(2) AIA(Gils, 2024). There are strong reasons to believe it is. In the latter part of the first sentence, Article 50(2) AIA mentions the obligation to mark the synthetic content as generated or manipulated. Recital 133 AIA explicitly also mentions the manipulation of existing inputs. Thus, we assume here that the generation and manipulation of input are covered by Article 50(2) AIA.
@@
In contrast to the core deep fake regulation in Article 50(4) AIA, here, text content is also listed. Thus, the obligation of Article 50(2) AIA includes deep fakes but also extends beyond core deep fake regulation.
@@
The main obligation is to mark the synthetic content in a machine-readable format to ensure that the content is detectable as artificially generated or manipulated. On the technical side, watermarking synthetic outputs is far from a trivial task (Gils, 2024), although Recital 133 ALA explicitly mentions some techniques. It remains to be seen whether the community will invest some joint effort to develop global standards (Romero Moreno, 2024).
@@
Article 50(2) AIA does not explicitly mention that the marking of the synthetic output has to be detectable by human viewers [^10] We specifically cover image content in our paper. per se. Thus, it remains unclear whether there shall be a human-visible declaration of the synthetic content (Gils, 2024). However, Article 50(5) AIA states that the information referred to in Article 50(2) shall be provided to natural persons in a clear and distinguishable manner. Marking content only in a machine-readable way without visibility to humans would be deceptive. So in practice, it does not matter whether Article 50(2) AIA alone or in conjunction with Article 50(5) AIA requires some human-visible marking. Additionally, providers of (large) social media platforms already have an obligation to label deep fakes through Article 35(1)(k) DSA (Block, 2024).
@@
In contrast to the aforementioned minor issues in the transparency obligation, two exceptions in Article 50(2) AIA pose a major difficulty. Article 50(2) AIA states that
@@
This obligation [of Article 50(2) AIA] shall not apply to the extent the AI systems perform an assistive function for standard editing or do not substantially alter the input data provided by the deployer[…]
@@
So the first exception targets systems that perform assistive functions for standard editing tasks. However, the intention of the legislator remains unclear (Block, 2024). There is no legal definition or a clarification in the recitals for either of the terms “assistive function” and “standard editing”. What characterizes standard editing given the existence of modern AI-based image editing tools, such as Samsung’s scene optimizer or Google’s Best Take? The threshold for an assistive function or a standard editing tool is not fixed.
@@
From our point of view, both exceptions [^11] There is also a third exception for law enforcement purposes which we do not consider in our paper. pose a serious challenge to existing image editing tools on smartphones and computers. We discuss this issue in depth in Section 4.
The other main transparency obligation regarding deep fakes is given in Article 50(4) AIA subparagraph 1. Here, deep fakes are specifically mentioned:
@@
In contrast to Article 50(2) AIA, this obligation only targets the deployers of an AI system. Furthermore, only deployers of systems processing image, audio, or video are regulated. Text systems, on the other hand, are not considered as deep fakes under the AI Act, and are regulated in Article 50(4) AIA subparagraph 2. It is unclear whether Article 50(4) AIA subparagraph 1 requires the AI content itself to be labeled—or whether labeling the system, for example, through a banner on the webpage, is sufficient. However, Recital 134 AIA clearly indicates that the AI content itself needs a label disclosing the artificial content. Thus, deployers of an AI system that create deep fakes need to ensure that humans can observe the origin (Romero Moreno, 2024).
@@
Compared to the non-trivial exceptions in Article 50(2) AIA, the exceptions for the deployers in Article 50(4) AIA are clearer. There are exceptions and limitations to this obligation for the law enforcement regime as well as for evidently artistic, creative, satirical, fictional or analogous work. The latter exception refers to the freedom of expression and artistic creation protected by Articles 11 and EU Charter of Fundamental Rights (Romero Moreno, 2024).
@@
The interplay between the obligations of the provider in Article 50(2) AIA and the obligations of the deployer in Article 50(4) AIA is complex and remains largely unclear. Recital 134 AIA explicitly mentions that further to the technical solutions employed by the providers of the AI system, the deployers also have to comply with the transparency obligations. It seems questionable whether double labeling adds a benefit to transparency. Standardized labeling efforts are needed (Gils, 2024).
@@
Both obligations try to regulate a technology that is, without the underlying intention of the user, neutral (Łabuz, 2024; De Ruiter, 2021). However, the overarching theme of both obligations is to ensure that even in cases where AI systems can create large amounts of synthetic images, humans are able to differentiate between synthetic and real images, see also Recital 133 AIA. This focus on the recipients of the synthetic content has already been criticized, as it does not focus on the victims of deep fakes, for example, in the case of (deep) pornographic images or videos (Veale and Borgesius, 2021; Łabuz, 2024; Meskys et al., 2020), which can cause severe harm to people. However, the non-compliance of providers or deployers of (commercial) AI systems with Article 50 AIA can be subject to a fine of up to €15,000,000, or, if the offender is an undertaking, up to 3% of its total worldwide annual turnover for the preceding financial year, whichever is higher (see Article 99(4)(g) AIA). Thus, non-compliance can be very costly for a company.
3.3. Summary of main critical legislative issues in Article 50 AI Act